40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 5 of 57  Not logged in ELOG logo
ID Date Author Type Category Subject
  2627   Fri Jul 30 17:28:14 2021 RadhikaDailyProgressCDSConnecting CTC100 to EPICS/rtcds system

[Koji, Radhika]

We were able to successfully integrate the Megastat CTC100 temp data into EPICS and Frame builder. Picking up from my last entry, we completed the following steps:

1. Restarted the EPICS st.cmd service - this "connected the wires" from the CTC100 to EPICS and allowed us to be able to read out live-time channel values. Koji created a script file to restart st.cmd:

cd /opt/rtcds/caltech/c4/target/qil-nfs/CTC100/iocBoot/iocCTC100

Note that if these channels need to be edited, this is the script one should call to restart EPICS. At this point, we could get the values by caget commands like


P.S. Chris sent us a permanent solution for the service. The process was killed and the service was started by the following command

systemctl restart CTC100

2. Appended new channel names to an existing .ini file (/opt/rtcds/caltech/c4/chans/daq/C0EDCU.ini). This allowed us to record these channels in Frame builder. I was originally going to create a new .ini file with these channels and add the name of the file to master, but we could not find the master file. Plus, instructions of fb4 told us to append to C0EDCU.ini specifically.

3. Restarted daqd@standiop.service on fb4:

sudo systemctl restart daqd@standiop.service

We used DataViewer to verify that the channels were being recorded. Success!

  2626   Fri Jul 30 14:44:56 2021 StephenThings to BuyCryo vacuum chamberUpgrades and updates in advance of PD testing

(Aidan, Stephen, Koji)

Building off of QIL/2597 after more thorough discussion in Mariner meeting today. Aidan and I will confirm details today and make moves toward installation of PDs next week.

Necessary capabilities (note: no need to scavange anything from the IRLabs dewar):

 - Cold temperatures = ready (via conductive mounting we have seen workpiece ~ 80 K, workpiece heating and CTC100 temp control is demonstrated)

 - Optical interface = ready (existing 1700 nm AR-coated window will be used at 2 micron, input power will just be calibrated with a power meter)

 - Electrical interface = almost ready, pending cabling (the cryo vacuum chamber has RTDs instrumented, so we just need the leads for the PDs, and we can create Cu twisted pair leads with in-vac crimp/solder sockets a la Koji's OSEM cabling) (also need in-air cabling with DB9 plug - trivial)

Improvements in advance of PD testing:

 - RTD cabling and Heater cabling (in vacuum) should be split with in-vac pins and insulated with PTFE tube - Koji has the magic material recommendations

Desired improvements (not necessarily in advance of PD testing):

 - Cu solid linkage (being fabricated).

 - Inner radiation shield should be clamped well to cold plate (consistent with most recent trial, but do we want better/more clamps?).

 - RTD mounting option on shields without cryo varnish (new threaded hole, new clamps).

 - RTDs consistent with planned Mariner RTDs (ref. QIL/2590).

 - Heater mounting should be directly to the workpiece via an improved clamp on the 2" x 2" grid (rather than on unused Si cantilever workpiece holder).

  2625   Fri Jul 30 12:22:56 2021 KojiSummaryCryo vacuum chamberCooling curve comparisons

In all aspects, the latest cooling shows the best performance thanks to better thermal connection, thermal isolation, and the black paint.

- The cold head cooling is faster and cooler

- The inner shield cooling is faster

- The test mass cooling is faster

Attachment 1: comparison_cold_head.pdf
Attachment 2: comparison_inner_shield.pdf
Attachment 3: comparison_test_mass.pdf
  2624   Fri Jul 30 12:22:15 2021 RadhikaDailyProgressCDSConnecting CTC100 to EPICS/rtcds system

Currently the CTC100 temperature data for Megastat is extracted via ethernet (telnet) through a python file. The file queries the device for temperature readings every minute and stores the data to a .txt file on QIL-WS1. Our goal is to connect the CTC100 data directly to SLOW EPICS in the QIL.

It has been helpful that the other CTC100 in the QIL (henceforth called the original CTC100), which is used for temp monitoring of the small PD testing chamber, had already been integrated into EPICS. I located the relevant files (from QIL-WS2) under /opt/rtcds/caltech/c4/target/qil-nfs/CTC100/. This will now be the called parent_directory.

The CTC100 protocol was already defined at parent_directory/iocBoot/iocCTC100/protocols/ctc100.proto. I added protocol functions corresponding to the commands I would need for querying the Megastat CTC100, of the form:

read_MS_WorkPiece {
    out "WorkPiece?";
    in "%f";
    @init {
      out "WorkPiece?";
      in "%f";

I created a function block for each of the 4 channels read by the Megastat CTC100.

Next, I located the st.cmd file: parent_directory/iocBoot/iocCTC100/st.cmd. I added the following line to configure the port to the new CTC100, copying the sytax of the original CTC100:

drvAsynSerialPortConfigure("CTCMS", "", 0, 0, 0)

The .db file for channels corresponding to the original CTC100 is found at parent directory/db/ctc100.db. I added blocks for the new EPICS channels to write the CTC100 data into:

    field (INP, "@ctc100.proto read_MS_WorkPiece CTCMS")

Here, the input field indicates the path to the protocol and which function in the protocol to call; and which port to communicate with. I added a block for each of the 4 channels of the Megastat CTC100.

The next step should be to restart the EPICS service so that these channels can be created. I have not been able to locate the right service file to restart, but hopefully once I do, I should be able to call caget on one of the channels and see a real-time value.

Next, I will create a .ini file to load into the frame builder service, so that the frame builder can record the new channels into frames and save records of data. I tried to look for this .service file with the help of Aidan and Anchal, but we have not yet been able to locate it. I hope to solicit Chris' help for both this task and the one above.


  2623   Fri Jul 30 09:48:11 2021 RadhikaDailyProgressCryo vacuum chamberCooling model updates from Koji's analysis

I replicated Koji's recent cooldown analysis on data prior to painting the test mass with black coating. The model first considers conductive cooling of the coldplate + inner shield from the cold head, via copper braid. Then it considers radiative cooling of the test mass from the coldplate + inner shield. 

To model the conductive cooling of the coldplate + inner shield, I used:

dT_coldplate/dt = C * k_Cu(T) * A_Cu/l_Cu * (T_coldhead - T_coldplate) / (Cp_Al(T) * m_coldplate)

where A_Cu and l_Cu are the cross-sectional area and length of the copper braid. I used a temperature-varying heat capacity of Cu and specific heat of Al. In order to align this model with the data, I found that the constant C=0.085. I am not sure what extra factors should be contributing to this scaling, but once it is added, the model aligns well with the two time constants apparent in the data [Attachment 1]. I will connect with Koji to determine what he considered here / if there is something I am missing.

I then took the aligned coldplate + inner shield cooling model to consider the radiative cooling of the test mass. I used:

dT_tm/dt = Fe * sigma * A_tm * (T_coldplate^4 - T_tm^4) / (Cp_Si * m_tm)

where we assume T_coldplate is the temperature of the coldplate + inner shield. Fe is the emissivity coefficient Koji considers in his analysis, which I found to be consistent with his result: 0.15. As he stated in a previous entry [2617], Fe can be broken down as:

1/Fe = 1/e_tm + (1/e_surr - 1)*A_tm/A_surr,

where e_surr and A_surr are the emissivity and area, respectively, of the surrounding coldplate and inner shield. Using e_surr=0.07 (rough Al), we get that an Fe value of 0.15 corresponds to a test mass emissivity of 0.18. This is slightly lower than Koji's value, due to differences in our calculated surface areas, but otherwise consistent.

A key point is that once the conductive cooling of the coldplate is modeled accurately (with a fudge factor of 0.085), the radiative cooling model of the test mass lines up well with the data without the need for another fudge factor. Note that the radiative cooling model above does not use a temperature-varying specific heat of Si [Attachment 1]. If a temperature-dependent value is used, we end up with the test mass cooldown model seen in Attachment 2. This causes the model to diverge from the data, so another factor might be missing in the model. Perhaps using temperature-dependent emissivities will correct for the deviation and cause even better agreement. This is a future step for the model.

Lastly, the painting of the test mass will increase its emissivity value, strengthening the radiative link between the test mass and its surroundings. (Koji has already posted updates on this cooling trend, and I will use this data once I obtain a copy.) Based on Koji's entry [2617], we can consider a new e_paint of 0.5 and 1. Attachment 3 compares radiative cooling models of the test mass using different emissivity values. We can expect that if the coating performs as expected, the test mass can reach 123K between ~87-110 hours. A next step is to plot the cooldown data for the painted test mass to see how accurate this prediction is. 

I will next aim to understand the 0.085 fudge factor needed to align the conductive cooling model with the coldplate cooling data. I will also add a fitting feature to directly spit out the optimal factors needed in both conductive and radiative cooling. 


Attachment 1: cooldown_Cp_fixed.pdf
Attachment 2: cooldown_Cp_varying.pdf
Attachment 3: e_tm_comp_Cp_fixed.pdf
  2622   Thu Jul 29 13:11:17 2021 KojiSummaryCryo vacuum chamberCooling progress: Update

The current cooling curve suggests that the radiative cooling factor Fe (black body =1) increased from 0.15 to 0.5.

Update: The test mass temp is reaching 200K at ~27hrs. cf previously it took 50hrs
Update: The test mass temp is 170K at ~41.5hrs.

OSEM illumination & photodetector efficiency has been kept increasing @41.5hrs

Attachment 1: temp_log_cool_down_20210728_1830.pdf
Attachment 2: cooling_model1.pdf
Attachment 3: cooling_model2.pdf
Attachment 4: OSEM_cooling.pdf
  2621   Thu Jul 29 00:42:38 2021 KojiSummaryCryo vacuum chamberThe test mass successfully suspended

[Stephen Koji]

Road to cooling down

  • The suspension with the test mass was installed in the chamber again
  • Looking at the oplev beam, we jiggled the wire loop position to adjust the alignment approximately.
  • The oplev beam was aligned more precisely.
  • We intentionally kept the OSEM at the "fully-open" position, while it is still close to the magnet so that we can have some actuation.
  • The coil driver was tested before closing the chamber, but it did not work.
    The coil itself was still intact, and the mirror was responding to the coil current if the coil current of ~100mA was applied from a bench power supply with the current ~100mA).
    So the problem was determined to be external.
  • Once we were satisfied with the oplev/OSEM conditions, the inner and outer lids were closed. Then the chamber was closed.
  •  Started pump down.
  • Started cooling down @18:30 / started temp logging too. Log filename: temp_log_cool_down_20210728_1830.txt

The photos were uploaded to Google Photo of WB labs.

The coil driver issue was resolved:

  • It was necessary to take care of the enable switch. Made a DB9 short plug for this purpose.
  • The output R was 1.2K (i.e. 2.4K across the + and - outputs). We needed ~10x more to see visible motion of the mass
  • e.g. The internal gain of the driver is x1.1. If we connect 5VDC input across the diff input of the driver yields, +11V shows up across the outputs of the final stage.
    If the R across the coil is ~100Ohm, we get ~100mA.
  • Soldered 6 x  330Ohm (1/8W) in parallel to 1.2K R_out. -> This ended up 51.5Ohm x2 across the coil. Each R=330 consumes ~1/10W. ->OK

Checking the DAQ setup / damping loop

  • DAQ setup
    • ADC: QPD X->FM16 / Y->FM17 / S->FM18 / OSEM-> FM19
    • DAC: CH11 -> Coil Driver In
  • Connected FM16 and FM17 to the coil drive by setting C4:TST-cdsMuxMatrix_12_17 and C4:TST-cdsMuxMatrix_12_18 to be 1.0
  • It was not obvious if the coil could damp the rigid body modes.
    • Actating the magnet caused Yaw motion most. Some Pitch motion too.
    • Configured FM16 and FM17 for the damping loop.
      • Filter Bank #1: [Diff0.1-10]  Zero 0.1Hz / Pole 10Hz
      • Filter Bank #10: [Anti Dewht]  Zero 1&200Hz / Pole 10&20Hz
    • Tried various damping gain. The mass was moving too much and the proper gain for the damping was not obvious.
    • So, the initial damping was obtained by shorting the coil at the coil in of the sat amp unit. (Induced current damping)
    • Once the test mas got quieter, it was found that -0.01 for FM16 could damp the yaw mode. Also it was found that +0.1 for FM17 could damp the pitch mode. (But not at once as the filters were not set properly)
  • TF measurement for calibration
    • The beam was aligned to the QPD
    • The test mass was damped by using the damping loops alternately 
    • Taken a swept sine measurement Filename: OSEM_TF_210729_0243.xml
      Recorded the time, saved the data, and took a screenshot
      • This measurement was taken @T_IS=252K / T_TM=268K @t=8hr (2:30AM), Rcoil=15.6Ohm
    • Second measurement Filename: OSEM_TF_210729_2147.xml
      • @T_IS=172K / T_TM = 201K @t=27.5hr (10PM), Rcoil=10Ohm
    • 3rd measurement Filename: OSEM_TF_210730_1733.xml
      • @T_IS=116K / T_TM = 161K @t=47hr (5:30PM), Rcoil=?
    • 4th measurement Filename: OSEM_TF_210731_2052.xml
      • @T_IS=72K / T_TM = 134K @t=75hr (9:30PM), Rcoil=6.0Ohm


  • The Satellite amp brought from the 40m is used as-is.
  • The initial OSEM reading was 8.8V, this corresponds to ~30000cnt.
  • As the OSEM was cooled, this number was increasing. To avoid the saturation, a voltage divider made of 4x 15kOhm was attached. I didn't expect to have the input impedance of the AA filter (10K each for the diff inputs), this voltage divider actually made 18.24V across POS and NEG output to be 5.212V to the AA fiter. So the voltage division gain is not 0.5 but 0.2859.
  • This made the ADC range saved, but we still have a risk of saturating the PD out. If this happens. The PD TIA gain will be reduced before warming up.
    -> The TIA and whitening stages use AD822, and the diff output stage uses AD8672. AD822 can drive almost close to rail-to-rail. AD8672 can drive upto ~+/-14V.

There was not enough time for the QPD calib -> Tomorrow

  2620   Wed Jul 28 00:59:47 2021 KojiSummaryCryo vacuum chamberThe test mass successfully suspended

[Stephen Koji]

While Stephen worked on the RTD reattachment, I worked on the suspension part.

- First of all, we found that the magnet was delaminated from the silicon mass (Attachment 1). It was bonded on the test mass again.

- The suspension frame was tweaked so that we have ~max suspension length allowed.

- The first attempt of suspending the mass with steel wires (0.0017" = 43um dia.) failed. Stephen and I went to downs and brought some reels.

- I chose the wire with a diameter of 0.0047" (= 119um) (Attachment 2). ~8x stronger! The suspension was successfully built and the mass is nicely sitting on the 4 strain releasing bars (improvised effort). (Attachments 3/4)

We can install the suspension in the chamber tomorrow (today, Wed)!


Attachment 1: P_20210727_154143.jpeg
Attachment 2: P_20210727_190356.jpeg
Attachment 3: P_20210727_190426.jpeg
Attachment 4: P_20210727_190543.jpeg
  2619   Mon Jul 26 22:49:00 2021 KojiSummaryCryo vacuum chamberAquadag painting

[Stephen Koji]

We decided to paint the silicon test mass with Aquadag to increase the emissivity of the test mass.

Stephen brought the Aquadag kit from Downs (ref. C2100169) (Attachment 1)

It's a black emulsion with viscosity like peanut butter. It is messy and smells like squid (Ammonium I think) (Attachment 2)

We first tried a scoop of Aquadag + 10 scoops of water. But this was too thin and was repelled easily by a Si wafer.
So we tried a thicker solution: a scoop of Aquadag + 4 scoops of water. (Attachment 3)

The thicker solution nicely stayed on the Si wafer (Attachment 4)

We want to leave the central area of the barrel unpainted so that we can put the suspension wire there without producing carbon powder. (Attachment 5)
1.5" from the edge were going to be painted. The central1" were masked.

The picture shows how the Si test mass was painted. The test mass was on a V-shaped part brought from the OMC lab. The faces were also painted leaving the mirror, while the place for RTD, and the magnet were not painted. (Attachment 6)

It looked messy while the painting was going, but once it started to dry, the coating looks smooth. It's not completely black, but graphite gray. (Attachment 7)

After the test mass got dry, another layer was added. (Attachment 8)

Then made it completely dry. Now the mask was removed. Nice! (Attachments 9/10)

Attachment 1: 20210726164254_IMG_0768.jpeg
Attachment 2: 20210726164530_IMG_0769.jpeg
Attachment 3: 20210726164225_IMG_0766.jpeg
Attachment 4: 20210726164957_IMG_0772.jpeg
Attachment 5: 20210726173608_IMG_0774.jpeg
Attachment 6: 20210726174523_IMG_0775.jpeg
Attachment 7: 20210726182715_IMG_0783.jpeg
Attachment 8: 20210726192042_IMG_0784.jpeg
Attachment 9: 20210726192837_IMG_0790.jpeg
Attachment 10: 20210726192853_IMG_0791.jpeg
  2618   Mon Jul 26 01:30:42 2021 KojiSummaryCryo vacuum chamberPrep for the 2nd cooling of the suspension

Updated Jul 26, 2022 - 22:00


  1. Reconstruct the cryostat
    1. [Done] Reinstall the cryo shields and the table (Better conductivity between the inner shield and the table)
    2. [Done] Reattach the RTDs (Inner Shield, Outer Shield)
      -> It'd be nice to have intermediate connectors (how about MIllMax spring loaded connectors? https://www.mill-max.com/)
    3. Reattach the RTD for the test mass
  2. Test mass & Suspension
    1. [Done] Test mass Aquadag painting (How messy is it? Is removal easy? All the surface? [QIL ELOG 2619]
    2. [Done] Suspension geometry change (Higher clamping point / narrower loop distance / narrower top wire clamp distance -> Lower Pend/Yaw/Pitch resonant freq)
    3. [Done] Setting up the suspension wires [QIL ELOG 2620]
    4. [Done] Suspend the mass
  3. Electronics (KA)
    1. [Done] Coil Driver / Sat Amp (Power Cable / Signal Cables)
    2. Circuit TF / Current Mon
    3. [Done] DAC wiring
    4. [Done] Damping loop
  4. Sensors & Calibration (KA)
    1. [Done] Check OSEM function
    2. [Done] Check Oplev again
    3. Check Oplev calibration
    4. [Done] Check Coil calibration
    5. Use of lens to increase the oplev range
    6. Recalibrate the oplev
  5. DAQ setup (KA)
    1. [Done] For continuous monitoring of OSEM/OPLEV
  2617   Sun Jul 25 21:45:46 2021 KojiSummaryCryo vacuum chamberAbout the radiation heat transfer model

The following radiation cooling model well explained the cooling curve of the test mass (until ~150K)

\dot{Q}=0.15 A\,\sigma (T_{\rm SH}^4-T_{\rm TM}^4)

where dQ/dt is the heat removed from the test mass, A is the surface area of the test mass, \sigma is the Stefan-Boltzmann constant, T_SH and T_TM are the temperatures of the surrounding shield and the test mass.

Can we extract any information from this "0.15"?

I borrowed "Cryogenic Heat Transfer (2nd Ed)" by Randall F. Barron and Gregory F. Nellis (2016) from the library.
P.442 Section 8.5 Radiant Exchange between Two Gray Surfaces can be expressed by Eq 8.44

\dot{Q} = F_e F_{1,2} \sigma A_1 (T_2^4-T_1^4)

where T_i is the temperature of objects 1 and 2. For us, OBJ1 is the test mass and OBJ2 is the shield. A1 is the surface area of A1. F_1,2 is the view factor and is unity if all the heat from the OBJ1 hits OBJ2. (It is the case for us.)

F_e is an emissivity factor.

The book explains some simple cases in P 443:

Case (a): If OBJ2 is much larger than OBJ1, F_e = e_1 where the e_i is the emissivity of OBJi. This means that the radiated heat from OBJ1 is absorbed or reflected by OBJ2. But this reflected heat does not come back to OBJ1. Therefore the radiative heat transfer does not depend on the emissivity of OBJ2.

Case (b): If OBJ1 and OBJ2 has the same area, \frac{1}{F_e} = \frac{1}{e_1} + \frac{1}{e_2} -1. The situation is symmetric and the emissivity factor is influenced by the worse emissivity between e1 and e2. (Understandable)

Case (c): For general surface are ratio,  \frac{1}{F_e} = \frac{1}{e_1} + \left(\frac{A_1}{A_2}\right)\left(\frac{1}{e_2} -1 \right ). OBJ2 receives the heat from OBJ1 and reradiates it. But only a part of the heat comes back to OBJ1. So the effect of e2 is diluted.

For our case, OBJ1 is the Si mass with DxH = 4in x 4in, while the shield is DxH = 444mm x 192mm. A1/A2 = 0.12.
We can solve this formula to be Fe=0.15. e1 = (0.147 e1)/(e2-0.0178).

Our inner shield has a matte aluminum surface and is expected to have an emissivity of ~0.07. This yields the emissivity of the Si test mass to be e1~0.2

How about the sensitivity of e1 on e2? d(e1)/ d(e2) = -0.95 (@e2=0.07).

Can Aquadag increase the radiative heat transfer?

Depending on the source, the emissivity of Aquadag varies from 0.5 to 1.
e.g. https://www.infrared-thermography.com/material-1.htm / https://www.mdpi.com/1996-1944/12/5/696/htm

  • Assuming Aquadag's emissivity is ~1
    • If only the test mass is painted, F_e increases from 0.15 to 0.39 (x2.6)
    • If the inner shield is also painted, F_e increases to 1, of course. (pure black body heat transfer)
    • If shield panels are placed near the test mass with the inner surface painted, again F_e is 1.
  • Assuming Aquadag's emissivity is ~0.5
    • If only the test mass is painted, F_e increases from 0.15 to 0.278
    • If the inner shield is also painted, F_e increases to 0.47.
    • If shield panels are placed near the test mass with the inner surface painted, F_e is 0.33 assuming the area ratio between the test mass and the shield panels to be unity.

It seems that painting Aquadag to the test mass is a fast, cheap, and good try.

  2616   Fri Jul 23 20:53:40 2021 KojiSummaryGeneralJul 17, 2021: Canon camera / small silver tripod / macro zoom lens / LED ring light returned / ELectronics borrowed

[Returned] Brought one HAM-A coil driver (D1100687 / S2100619) and one Satellite Amplifier (D1002818 / S2100741) from the 40m

Also brought some power cables.

Brought ~1m of 0.0017" (~43um) misical wire. This will make the tension stress be 341MPa. The safety factor will be ~7.


Attachment 1: P_20210723_212158.jpg
  2615   Thu Jul 22 22:03:45 2021 KojiSummaryCryo vacuum chamberTest mass heating in progress (2021/07/21 ~ 2021/07/23)

- Temperature Log updated 2021/7/23 12:00 Heating Ended

- Assuming reaching the room temp at ~30hrs and heating power saturated at 22W: Predicted heat injection 30*3600*22 = ~2.4MJ

Update from Stephen
- Note that we can check logging accuracy against the snapshot (timestamp 20210723_1113).
If my math is correct, this would be time = 37.35 38.35 hours

Update from KA
=> The corresponding time in sec is 138060 sec
The raw data line for the corresponding time is:

138016.839614, 295.805, 306.678, 302.518, 312.401, 0.000, 0.000, -0.001, 0.621, 0.622, 1.429, 0, 0, NaN, NaN, NaN
The values on the photo 295.806, 306.677, 302.518, 312.401 ==> Well matched. Victory!

Attachment 1: IMG-9395.jpg
Attachment 2: temp_log_warmup_20210721_2052.pdf
  2614   Wed Jul 21 21:05:59 2021 KojiSummaryCryo vacuum chamberTest mass cooling (2021/07/16 ~ 2021/07/21)

[Stephen and Koji for discussion / Koji for the execution]

1. Temperature Trend

See [QIL ELOG 2611] for the updated temp log and the cooling model.

Considerations for the next cycle:
-> How can we accelerate the cooling? It seems that the table cooling is conduction limited. Improve the cold head connection.
-> We want to move the RDTs
-> How can we improve radiative cooling?

2. Oplev Trend (Attachment 1)

Sum: The beam has been always on the QPD (good). See also Attachment 2

X&Y: In the first few hours the beam drifted in -X and then +X while Y had slow continuous drift in +Y. ~11hours later sudden drift in -Y and totally saturated. Also -X saturation observed @~16hrs. Again +Y drift was seen @~25hrs. The totally saturated in -X and +Y.
They may be related to the drift of various components with various cooling time scale.

Visual check: ~2mm shift in X&Y is visually observed. Attachment 2

-> How can we quantify the drift? What information do we want to extract?

3. OSEM and the magnet

The magnet is intact. And the suspension seemed still free after cooling (Attachment 3)
Significant misalignment was not visible. No visible damage by cooling was found. The coil is alive and the PD/LED are also intact. Fluke showed that they are still diodes, but their function was not checked.

The coil resistance changed from 16Ohm -> 4.2Ohm. For the 16Ohm, 2 Ohm was from the wire. Let's assume we still have 2Ohm overhead -> The coil R changed from 14->2.2. This corresponds to the coil temperature of the order of ~100K. This is not so crazy.

Some actuation current was applied to the magnet. For this test, the oplev was realigned.
First, some ~300mA current pulses were applied to the coil. The ringdown of the yaw mode was visible. Then the DC current of 100mA was applied. This didn't make visible change on the spot position but the data showed that there was a DC shift.

-> We prefer to have a softer suspension for the next test.

4. CTC100 logging

During the cooling we kept having inaccurate data logged compared with the displayed data on the screen of CTC100.
As soon as the cooling logging was stopped, telneting to CTC100 was available. So, I telnetted to the device and sent the data transfer command ("getOutput"). Surprisingly, the returned values agreed with the displayed values.
So my hypothesis is that somehow the data strings are buffered somewhere and gradually the returned values get delayed. From the behavior of the device, I imagined that the fresh telnet connection gives us the latest data and there is no buffering issue.

So I tweaked the data logging code to establish the telnet connection every time the values are asked. The connection is closed after the every data acquisition. I like this as we can also make the test connection between each data acquisition points, although I have not tried it yet. The code is in the same folder named ctc100_controller_v2.py

5. Heating

Now I thought that I did all I wanted to do this evening, so the heater was turned on at ~20:50, Jul 21. The heating power saturated at 22W, which is the set limit.

Attachment 1: oplev_trend.png
Attachment 2: 20210721201333_IMG_0765.jpeg
Attachment 3: 20210716234113_IMG_0742.jpeg
Attachment 4: Screenshot_from_2021-07-21_20-19-09.png
  2613   Wed Jul 21 14:53:28 2021 KojiSummaryGeneralJul 17, 2021: Canon camera / small silver tripod / macro zoom lens / LED ring light borrowed -> QIL

See https://nodus.ligo.caltech.edu:8081/40m/16250

  2612   Wed Jul 21 13:14:11 2021 PacoLab Infrastructure2micronLasersBrimrose AOM and amplifier

Yesterday I came in the QIL and performed an express kidnapping of the 2um in fiber AOM (Brimrose) and the 5 W RF amplifier that was hooked to the RF in port (though it seems it saturates at ~ 600 mW from past elogs). I will test it with the 1419 nm ECDL fiber pickoff port to see that it works and if it doesn't I will reinstall it in the 2um testing facility.

  2611   Tue Jul 20 17:28:30 2021 KojiSummaryCryo vacuum chamberA cooling model (Temp Log 210716_2255)

Updated the model the latest log data with cooling prediction

  • The radiative cooling is expected to be the dominant cooling mode.
  • It will take ~3 more days to reach 123K. We don't need to wait for it.
  • For more informative temp data, we need the temperature of the inner shield and the table.

  • We know the cold head temp from the measurement. For the prediction, the constant cold head temp of 65K was assumed.
  • The table temp was estimated using conductive cooling model + linear empirical dependence of the conductivity on the temp
  • The constant specific heat of the silicon mass (0.71 J/K/g) was assumed. This may need to be updated.
  • The radiative cooling is given from Stefan–Boltzmann law with the emissivity of 0.15 for both the shield and the mass.
  • The conductive cooling of the test mass was estimated using: Wire diameter 0.017" (=0.43mm), 4 wires, length of ~10cm (guess), no thermal resistance at the clamps (-> upper limit of the conductive cooling)

Radiative cooling already gives us a good agreement with the measured temp evolution for the test mass. The conductive cooling is not significant and does not change the prediction.

Updated the plot with the new data (2021/7/21 12:30PM)

Attachment 1: cooling_model.pdf
  2610   Tue Jul 20 11:33:52 2021 KojiSummaryCryo vacuum chamberA cooling model (Temp Log 210716_2255)

A naive cooling model was applied to the cooling curve.
A wild guess:

- The table temp is the same as the test piece temp as measured on 2021/7/9
- The inner shield temp is well represented by the table temp
- The specific heat of Si is almost constant (0.71 [J/(g K)] between 300K~200K

Radiative cooling:
The curve was hand-fitted by changing the emissivity of the inner shield and the silicon mass. I ended up having the same values for these to be 0.15.
Surprisingly well fitted!

Conductive cooling:
The conductive cooling through the wire does not fit the cooling curve, although the quantitative evaluation of the wire conductivity needs to be checked carefully.

Stephen shared attachments 2 and 3, which contain insights on the wire used to hang the Si mass. .017" diameter Music Wire from California Fine Wire, 2004 vintage, borrowed from Downs High Bay.

Attachment 1: cooling_model.pdf
Attachment 2: IMG_9390.JPG
Attachment 3: IMG_9391.JPG
  2609   Mon Jul 19 17:21:19 2021 KojiSummaryCryo vacuum chamberTemp Log 210716_2255

Temp Log on Jul 19 2021 17:20

I wonder what is the heat transfer mode for the test mass right now. Radiative? or Conductive through the wires?


Attachment 1: temp_log_cool_down_20210716_2255.pdf
  2608   Mon Jul 19 15:57:17 2021 StephenSummaryCryo vacuum chamberTemp Log 210716_2255

Uh oh, review of the cooldown plot from the previous cooldown (QIL/2603) shows workpiece temperature of ~92 K at conclusion, while a temperature of 65K was observed in the CTC100 readout (Attachment). The logging of the warmup is consistent with the CTC100 image, as the logging started a few minutes after the warmup was started, and the warmup "5 minutes after starting" temperature of ~ 71 K is a practical temperature.

Seems to be something weird going on here, we will need to have Radhika take a look on her return (and continue taking photos of the CTC100 whenever we stop by).


Temperature log for the first 2 hours (Attachment 1)

I wonder why the temperatures displayed on CTC100 and the ones logged are different...?



  2606   Sat Jul 17 00:55:41 2021 KojiSummaryCryo vacuum chamberTemp Log 210716_2255

Temperature log for the first 2 hours (Attachment 1)

I wonder why the temperatures displayed on CTC100 and the ones logged are different...?


Attachment 1: temp_log_cool_down_20210716_2255.pdf
  2605   Fri Jul 16 23:28:24 2021 KojiSummaryCryo vacuum chamberSus Test Work 07/16/2021

[Stephen Koji]

We started cooling down of the test mass.


- Stephen vented the chamber at 2PM. An optical port was moved to see the OSEM from the back.

OSEM wiring

- Brought DSub crimp sockets from the 40m. We picked up 3x 1m LakeShore WCT-RB-34-50 (twisted silver-plated copper, 34 AWG with Teflon insulation). The ends of the wires were dangled so that crimping is possible. A single wire resistance was measured to be ~1Ohm at room temp. (Attachment 1)

- OSEM pin out / backside view (cable going down) (Attachment 2)

|   o   o   o |
| o   o   o   |                 Wire
  ^ ^ ^ ^ ^ ^---PD K        ---- R3
  | | | | |-----PD A        ---- B3
  | | | |-------LED A       ---- B2
  | | |---------LED K       ---- R2
  | |-----------Coil End    ---- B1
  |-------------Coil Start  ---- R1

Twisted Pair 1: (R1&B1) with 1 knot  at the feedthru side
Twisted Pair 2: (R2&B2) with 1 knot  at the feedthru side
Twisted Pair 3: (R3&B3) with 1 knot  at the feedthru side

Dsub feedthru in-air pinout (Mating side)

    1  2  3  4  5
\ o  o  o  o  o /
 \ o  o  o  o  /
   6  7  8  9

Pin1 - Coil Start
Pin6 - Coil End
Pin2 - LED K
Pin7 - LED A
Pin3 - PD A
Pin8 - PD K

Pin1-6 R=16Ohm
Pin2-7 Diode V (with Fluke) 1.18V (Pin2 black probe / Pin7 red probe)
Pin3-8 Diode V (with Fluke) 0.7V (Pin3 red probe / Pin8 black probe)

- OSEM pin out / backside view (cable going down)

Suspension installation (Attachment 3)

- The sus frame was moved into the chamber

- We measured the test mass dimension before installation: L 3.977" D 4.054"

- The attached mirror size is 1"x1" made of SUS #8 (?)

- The mass was suspended. The height / rotation of the mass was adjusted so that the reflecting mirror is visible from the oplev window and also the OSEM magnet is visible from the OSEM window.

- The OSEM was placed on an improvised holder. (Attachment 4)

Oplev installation

- ...Just the usual oplev installation. Adjusted the alignment and the return beam hits right next to the laser aperture. This beam was picked off by a mirror and steered into a QPD. (Attachments 5/6)

- The lever arm length is ~38" (960mm) -- 9" internal / 29" external
- The oplev signal is shaking so much and occupying ~50% of the full scale. Added a lens with f=250 to make the beam bigger, but the improvement was limited.

Pumping down

- Started ~8:30PM?

DAQ setup

- Wired 3 BNC cables from the table to the DAQ rack. CHX/Y/S are connected to ADC16/1718ch.

- The real-time processes seemed dead. Looked at [QIL ELOG 2546] to bring them up. TIM/DAQ error remains, but the data stream seems alive now. Leave it as it is.


- Temp Logging started. Filename: temp_log_cool_down_20210716_2255.txt

- Cryocooler turned on. ~10:55PM

- Confirmed the cold head temp was going down. The cold head temp is 75K at 0:30AM

OSEM photo

- An example photo was taken from the rear window. The attempt with 40m's Canon failed. Attachment 7 was taken with KA's personal compact camera with a smartphone LED torch. The gap between magnet and OSEM is highly dependent on the view axis. So this is just a reference for now.

Attachment 1: 20210716170727_IMG_0719.jpeg
Attachment 2: 20210716174712_IMG_0723.jpeg
Attachment 3: 20210716195953_IMG_0726.jpeg
Attachment 4: 20210716200005_IMG_0728.jpeg
Attachment 5: 20210716200224_IMG_0734.jpeg
Attachment 6: 20210716200112_IMG_0733.jpeg
Attachment 7: 20210716234113_IMG_0742.jpeg
  2604   Thu Jul 15 23:37:53 2021 KojiSummaryCryo vacuum chamberBonding work for the prep of the preliminary suspension test

[Stephen / Koji]

Bonding work for the prep of the preliminary suspension test

- 1" sq mirror-ish polished SUS piece was bonded to a face of the silicon mass. We chose the location right next to a line on the barrel. (Attachment 1)

- The mass was flipped with two more same thickness pieces used for the spacers to keep the mass horizontal.

- A pair of an OSEM and dumbbell-magnet was brought from the 40m (courtesy by Yehonathan). The magnet was glued on the mass at the opposite position of the attached mirror because the optical ports are going to be arranged to share an axis. A piece of cryo varnish was also painted with a piece of cigarette paper at the center of the mass so that we can attach an RTD. (Attachment 2)

Next Things To Do (Attachment 3)

  • Vent the chamber
  • We will move an optical port to the opposite position of the other port.
  • A DB9 feedthru is going to be installed.
  • Suspension
    • Move the sus frame in the chamber
    • Suspend the mass
  • Sensor arrangement
    • Set up the oplev
    • Hold the OSEM at the height of the magnet
    • Set up a camera to observe the magnet-OSEM clearance
    • We improvise the DB crimping sockets so that we can electrically connect the OSEM (optional)
  • Pump down / cool down the chamber
    • The main target of the cooling is to check the cooling capability of the test mass mainly with radiative cooling.
    • An optional target is to observe the misalignment as a function of the temperature -
      • -> Oplev signals are to be connected to CDS / check if CDS is logging the data
    • Check if the OSEM/magnets survive the thermal cycle
    • If possible we can try to actuate the OSEM / check the LED/PD function at the cryo temp
Attachment 1: P_20210715_170102-1.jpg
Attachment 2: P_20210715_172218-1.jpg
Attachment 3: experiment_plan.pdf
  2603   Thu Jul 15 23:34:17 2021 KojiSummaryTempCtrlTemprerature Log for cooling down / warming up

[Stephen Koji Radhika]

Stephen and Radhika worked on the cooling down and warming up of the cryostat with the cold head RTD attached using a spring-loaded screw. No other configuration changes compared to QIL/2599. Here are the temperature log plots. Photos of spring clamped RTD are outstanding, but the clamp is the same as the workpiece pictured in QIL/2599/Attachment 12.

Attachment 1: temp_log_cooldown_20210709_1747.pdf
Attachment 2: temp_log_warmup_20210712_1315.pdf
  2602   Mon Jul 12 14:42:42 2021 StephenDailyProgressCryo vacuum chamberRTD attached to coldhead with spring clamp, Si mass to be installed this week

Brief summaries of the last week's progress and the coming week's plans (plots will be posted soon!):

- progress Friday 09 July: Opened the cryostat up at the cold head, and attached an RTD to the cold head with a spring clamp (instead of relying purely on the cryo varnish).

- progress Monday 12 July: Found 65 K workpiece temp and 63 K cold head temp. RTD was apparently held successfully by spring clamp, and we will continue to collect cold head temperature in future runs. Warmup was started, with old data collection completed (cooldown_20210709) and new data collection commenced (warmup_2021_07_12). Note that warmup started at 1:14 pm, and it took me ~ 5 minutes to stop and restart the script to changeover to the warmup data collection.

- table plan Wednesday 15 July: Complete in-air optical layout. Make one flat face of Si mass reflective.

- chamber plan Thursday 16 July: Open up main volume and drop in frame with Si mass. Connect RTDs. Start cooldown. Confirm cooldown is going ok (optical alignment, especially), and revert if necessary before things get too cold.

- table plan Friday 16 July: Maybe measure stuff, maybe better to wait till coming week and use controlled heating to hit different temperature setpoints.


  2601   Fri Jul 9 13:44:39 2021 RadhikaSummaryCryo vacuum chamber1D cooling model updates

*Takeaway*: The current 1D cooling model is getting closer to matching our observed cooling trends, mainly in the lower temperature limit. The predicted time constant is still much smaller than we are seeing in reality (by about a factor of 3), but this can potentially be improved by revising specific heat values and/or dimensional estimates for chamber components.

The model uses the known cooling power of the cold head [attachment 2] and considers radiative heat from the outer shield, baseplate (bottom lid), and mylar wrapping around braid. I increased the complexity of the script by solving a system of ODEs (for braid and coldplate temperature) simultaneously instead of assuming the temperatures are equal at all times, and solving only 1 ODE. This resulted in the model's lower temperature limit prediction matching our observed data, at ~66 K. 

The model still predicts a much smaller time constant than we are seeing. This is affected by specific heat values for Cu and Al, along with dimensional estimates of the coldplate and braid (AKA how much mass is being cooled). It is possible that these values are being underestimated in the model, which would lead to the smaller time constant. Currently the model uses constant values for the specific heat of Cu and Al (room temperature). But since specific heat increases with temperature, accounting for temperature dependence would lower the specific heat values and shift the model in the opposite direction (towards an even smaller time constant). Therefore I suspect the model is underestimating the mass of the coldplate, though I am unsure if this would completely correct the discrepancy. 

If the term (specific_heat * density * volume) of the coldplate (Al) is increased by a factor of 4, the model resembles the data well [attachment 3]. 

Attachment 1: model_vs_data_7-1.pdf
Attachment 2: coldhead_capacity.pdf
Attachment 3: model_vs_data_7-1.pdf
  2600   Fri Jul 9 10:57:58 2021 RadhikaSummaryCryo vacuum chamberCTC100 temperature extraction

I wrote a python script to extract temperature data from the CTC100 via ethernet, for monitoring cooldown/warmup of Megastat. This is intended to replace USB data extraction, which requires the user to manually insert/remove the stick and plug into a computer.

The script queries the CTC100 every ~60 seconds for the latest temperature values (the frequency can be supplied as a parameter, but default is 60s). The script writes line-by-line to a .txt file and also plot the outputted data once collection is terminated.

Here is a gitlab link to the script: https://git.ligo.org/voyager/mariner40/-/blob/master/CryoEngineering/ctc100_controller.py. It is also found on the QIL workstation at /home/controls/CTC100/ctc100_controller.py. To run from the workstation, open terminal to /home/controls (home). Then:

cd CTC100
python ctc100_controller.py --filename='tutorial'

 Here, 'tutorial' stands in for the desired filename for the outputted data. The script will start pulling data and will print each line to the terminal. It will continue printing and logging the temperature values until the user hits Ctrl+C in the terminal. This will terminate the script and output the final data file. The file is saved as a .txt file in /home/controls/CTC100/data.

Attachment 1: terminal1.png
  2599   Tue Jul 6 18:07:20 2021 RadhikaDailyProgressCryo vacuum chamberCooldown 7/1-7/6

I've attached cooldown data from 7/1-7/4, which was extracted via ethernet from the CTC100. Initially I was having issues getting the curves smooth; the data seemed very choppy and I assumed it was a device or ethernet issue. Turns out the csv file outputted by my script would cast all floats to integers if I opened it and saved, which was very strange. I'm still not sure why this happens, but my workaround is to not edit the csv at all after it is outputted.

The workpiece temperature stabilized at ~65.8 K, just below 66 K (last cooldown). This makes sense given the few minor tweaks we made in between (removed shorting Al foil, adjusted mylar / peek sheet insulation). The time constant is also consistent (~36.4 hrs). 

Unfortunately the RTD on the cold head appears to have detached yet again. It seems to have happened around the same time as the last cooldown. Stephen suggested replacing the current varnish bonding with a spring clamp, like what holds the workpiece RTD in place. 

I switched off the cryo cooler at 12pm today (Tue), so the chamber should be ready for venting Thursday or Friday. The tentative agenda is to add the spring clamp to secure the cold head RTD and cool down again (we will discuss any additional tweaks). I think I have an idea of why the current model predicts a much smaller time constant than we are seeing - I plan to work on this more this week. Hopefully once the cold head RTD stays fixed for an entire cooldown, this data can help tweak the model as well.

Attachment 1: cooldown_7-1.pdf
  2598   Fri Jul 2 11:27:01 2021 RadhikaDailyProgressCryo vacuum chamberCcooling updates 6/29-7/2

[Radhika, Stephen]

On Tuesday 6/29, we opened up Megastat after a weekend of warmup. Our goals were to 1) check on the RTD attached(?) to the cold plate to confirm that it had de-bonded, and 2) to check on the foil and mylar inside the chamber and around the copper braid. 

Indeed the tip of the cold head RTD had detached [attachment 1], but the mylar "cap" I created for insulation held up well. Inside the chamber, we noticed some shorting between aluminum foil on the inside of the outer lid and the outside of the inner lid [attachment 2]. Also, there was a gap in mylar around the copper braid, so some of the braid was visible. We also noticed that the strip of peek sheet inside the opening of the inner shield had slipped out of place, exposing the copper braid to the aluminum [attachment 8]. 

We rebonded the loose RTD to the cold head [attachment 3] and re-inserted the mylar "cap" [attachment 4 and 10]. Stephen suggested we add an additional piece of mylar to cover the portion of the cold head directly above the braid [attachment 5]. 

To prevent shorting inside the chamber, we trimmed the excess aluminum foil wrapped around the non-coated sides of the inner and outer shields [attachment 6]. We adjusted the mylar sheets around the copper braid to cover the previous gap, and used a mirror to ensure the bottom of the braid was covered [attachment 7]. We also realigned the peek sheet strip with the opening of the inner shield [attachment 9]. 

We reattached the heater to the workpiece [attachment 11] and the workpiece RTD with the spring clamp [attachment 12]. We placed back the inner and outer shield lids [attachment 13-14], closed up the chamber, and pumped down. Since I have been working on remotely extracting temperature data from the CTC-100, we chose to wait until the script was ready before turning on the cryocooler.

On Thursday 7/1, I got the temperature extraction script ready and turned on the cryocooler at ~5:15 PM. Fun addition: attachment 15 shows successful communication to the CTC-100 (I sent a command to display "hello" on the screen :D)

On Friday 7/2, I plotted the temperature trends so far [attachment 16]. As of 12pm Friday, the workpiece is at 177K. We will let the chamber cool over the weekend and see where we are on Tuesday.


Attachment 1: IMG_9126.jpeg
Attachment 2: IMG_9136.jpeg
Attachment 3: IMG_9147.jpeg
Attachment 4: IMG_9148.jpeg
Attachment 5: IMG_9145.jpeg
Attachment 6: IMG_9144.jpeg
Attachment 7: IMG_9142.jpeg
Attachment 8: IMG_9139.jpeg
Attachment 9: IMG_9152.jpeg
Attachment 10: IMG_9149.jpeg
Attachment 11: IMG_9153.jpeg
Attachment 12: IMG_9151.jpeg
Attachment 13: IMG_9154.jpeg
Attachment 14: IMG_9156.jpeg
Attachment 15: IMG_9158.jpeg
Attachment 16: cooldown_7_1-7_2.pdf
  2597   Tue Jun 29 17:29:16 2021 StephenThings to Buy2um PhotodiodesIntegrating 2um PD measurements into Cryo Vacuum Chamber

Radhika and I started discussing in detail the items needed to transfer the 2um PD setup into the Cryo Vacuum Chamber.

1) Electrical

To interface with the PD setup, it seems that we will need want to use the feedthrough already in use on the IRLabs cryostat. This square-flanged feedthrough appears to host the in-vacuum cabling. The in-air cabling would ideally also be reused.

To implement this, we will need to modify a blank 2.75" conflat flange (in hand in QIL boxes on wire shelf) with a the following features:

  • through bore, 1.18 diameter on center.
  • a 4x #4-40 tapped hole square bolt pattern, 1.062" spacing on center.
  • preferably, a hand-polished o-ring sealing area.

I will take care of this part through the PMA shop.

2) Optical

The Cryo Vacuum Chamber windows are Thorlabs VPCHW42‐C (ref. D2000310-v1). The broadband AR coating is rated for 1050-1700 nm (here's the data sheet), with a steep climb in AR coating reflectance. There is also fused silica substrate transmission to worry about. Perhaps the easiest way to proceed would be to repeat Koji's Transmission test (ref. QIL/2458) for the windows already in use on the Cryo Vacuum Chamber, and see if we can accept ~10% loss (or more).

Else, perhaps we scavange the IRLabs cryostat and create another custom blank CF modification?

Or buy a new, suitable window with CF interface? (<---- preferred option)

  2596   Tue Jun 29 17:16:29 2021 Ian MacMillanDailyProgressCryo vacuum chamberTemp Controller

Radhika and I were able to connect to the temperature controller using telnet. We simply reassigned the IP address of the temp controller to fit in the IP address range of the router/network. i.e. if the router is then the temp controller might be This shows that the router and network are working correctly again.

  2595   Tue Jun 29 15:45:43 2021 Ian MacMillanDailyProgressGeneralNetwork Issue

I was able to connect to the router (the one on top of the switch in the rack next to the orange tool drawers) and reconfigure it using the same internal IP as found labeled and mentioned in [2201]. The router now has a different external IP that I wrote on the router (twice). The network/router and switch is now working in the lab.

Update: It turns out that the router reset itself somehow. so any settings/ port forwarding that was set up before has probably been wiped out and would need to be set up again.

SSHing into the QIL needs to be set up again. 

  2594   Fri Jun 25 09:47:33 2021 RadhikaDailyProgressCryo vacuum chamberCooldown data

I've attached cooldown data from 6/21-6/24. Unexplicably, the CTC-100 stopped logging temperatures at ~50 hours after the start of cooling, or around 5pm Wednesday. The red logging indicator was definitely on at 3pm Wednesday, but it was off when I checked in Thursday afternoon. It must have somehow been disabled in between. On the bright side, the temperature of the workpiece stabilized at 66K. I had checked in at 53 hrs, 62 hrs, and 64 hrs and noted the temperatures by eye (green data points on attachment 2). 

The time constant for the workpiece is ~ 36.4 hrs, about the same as the previous run (tau = -40/ln(100/300)). While no improvements were made on this front (cooldown still very slow), it stabilized 6K below the previous run. 

The RTD at the cold head recorded a shoot-up in temperature right at the beginning of cooldown, which makes Stephen/me think that it debonded from the cold head. This temp reading would have been informative for our model, but we will verify it's status when we open up. We can find a better bonding mechanism for the next cooldown.

[Update] I let the chamber warm up over the weekend, and brought it up to room pressure on Monday 6/28. Next steps are to open up and investigate the cold head RTD and make improvements to bonding mechanism, if necessary. We will assess any other improvements (model driven) before the next cooldown.

Attachment 1: cooldown_6-21.pdf
Attachment 2: cooldown_model_vs_data_v2.pdf
  2593   Wed Jun 23 16:25:49 2021 Ian MacMillanDailyProgressGeneralNetwork Issue

I accidentally powered down and restarted the TP-link safe stream router and now nothing will connect to it. Not even things that are plugged into it. I can't even connect to the router's setup screen. I have no idea how to solve this issue short of resetting the router. I don't know how it was configured before so I will not reset it because I don't know how it was configured.

  2592   Wed Jun 23 15:03:12 2021 Ian MacMillanDailyProgressCDSTesting DAC2ADC code

I brought DAC2ADC_test.py code from the 40m to test in the QIL. I added a parser arg for the matrix root name (see attached code). I am running into the same problem as I did at the 40m where the channels seem to be locked to their values. I attached (attachment 2) the results text file that shows that for all inputs the outputs are the same. In the GDS screen the ADC is in red but I'm not sure how it got there or how to fix it.

run the test with these commands in the QIL

 $  cd IanMacMillan
 $  python DAC2ADC_Test.py -e 'C4:TST' -m 'cdsMuxMatrix'

One improvement that we could make to the code is have it zero out all of the matrix elements before it starts.

Attachment 1: QIL_CDS_TEST.tar.gz
  2591   Mon Jun 21 10:59:41 2021 RadhikaDailyProgressCryo vacuum chamberChamber updates 6/16-6/18

[Radhika, Stephen]

Background: There are currently 4 available RTDs for the large cryo chamber (henceforth named Megastat). We originally had 1 for the workpiece, 1 for the heater, 1 for the inner shield, and 1 for the outer shield. During the last sessions with the chamber open, the RTDs were moved around. Now there is 1 for the workpiece, 1 for the baseplate (bottom lid), one for the outer shield, and 1 for the cold head. These locations are marked in the diagram (attachment 9). 

On Wednesday 6/16, we opened up the chamber and removed the shields and coldplate. We added foil to the inner surface of the bottom lid (attachment 10) and attached an RTD there (we forgot to take a picture). Stephen made a new foil collar, and we decided to push it against the chamber walls as a better alternative to having the foil touch the coldplate (attachment 6). We added an extra layer of aluminized mylar to wrap the copper braid, and we fed another RTD through the braid tube to be attached to the coldhead. We lastly placed the coldplate and shields back in place.

On Friday 6/18, I attached the remaining RTDs (coldhead, outer shield, workpiece). We decided to remove the aluminum foil previously covering the coldhead, due to of fear of shorting to the tube wall. I taped several pieces of mylar together to cover the coldhead and insulate it from the wall (attachments 1-3). I placed the mylar contraption into the T (attachment 4) and then closed the bottom flange. I placed back the shield lids and the main lid of the chamber (attachments 7-8). I tried to pump down, but the pressure was stabilizing on the order of 1e-1 torr. Unsure of why the pressure wasn't decreasing, I turned off the pump and left for the day.

Today 6/21, I touched base with Stephen and we realized I forgot to replace the copper gasket in the bottom flange of the T. I then unscrewed and replaced the gasket. I pumped down and the pressure reached on the order of 1e-4 torr, so I proceeded with cooling. After a few hours I could see that the cold head RTD was reading a temperature around 180K. We will extract and plot cooling data late tomorrow or Wednesday.

Lastly, a chamber diagram is attached (attachment 9). The 4 RTD locations are marked by a red 'x'. The chamber components are numbered in blue (detailed below):

1. Workpiece

2. Heater

3. Inner shield

4. Outer shield

5. Copper braid (wrapped with mylar)

6. Cold head

7. Coldplate

8. Baseplate (bottom lid)

Attachment 1: IMG_1735.jpeg
Attachment 2: IMG_1738.jpeg
Attachment 3: IMG_1737.jpeg
Attachment 4: Photo_on_6-18-21_at_15.36.jpeg
Attachment 5: Photo_on_6-18-21_at_16.32.jpeg
Attachment 6: IMG_8947.jpeg
Attachment 7: Photo_on_6-18-21_at_16.33.jpeg
Attachment 8: Photo_on_6-18-21_at_16.33_#2.jpeg
Attachment 9: New_Note.png
Attachment 10: IMG_8973.jpeg
  2590   Fri Jun 18 10:15:14 2021 StephenUpdateCryo vacuum chamberTemperature sensor considerations

RTD thoughts - we have just been using the sensors that were provided, without noticing their constraints or deficiencies.

  • our RTD p/n is 615-1123-ND (Digikey), which is a Littlefuse PPG102A6 platinum RTD and has an apparent temperature range of -200 °C (73 K) to + 600 °C. The primary data sheet does not have a Resistance vs Temperature curve, simply presenting the slope of the temperature dependence in a parameter "Temperature Coefficient of Resistance" (TCR), but a table is available in an auxiliary document called "RT Chart". An image of this chart is attached below.
  • we have not violated that 73 K lower limit yet, but we are about to as we mount an RTD directly to the coldhead. Let's see how that RTD on the coldhead will withstand the lower temperatures.
  • the CTC100 manual indicates that any arbitrary calibration curve may be input, but there are a number of sensors' calibration curves built in - it might be a good idea to make our next sensor decision based on that list.
  • there is a Lakeshore guide and category page for temperature sensor selection - reconfiguring our 4 RTDs would cost over a kilobuck through Lakeshore, but perhaps we can learn general ideas from the guide as well.
    • it seems that switching to their Cernox line would be helpful in terms of packaging options, and would be the most accurate.
    • their line of silicon diodes would be suitable and has flexible packaging as well.
    • their platinum RTDs also have low temperature range and would be suitable. Packaging is cylindrical, so might be best to pursue the aluminum housing with a bolt hole.

Planning for next steps:

  • for now, it seems that we could get by with our generic and 73 K limited RTDs, and this option is tempting as it requires no additional effort.
  • if we decide we really want to have reliable sensing down to sub-50 K temperatures, we should move to one of the Lakeshore product lines (hopefully one which the CTC100 is configured with a calibration curve for) for about a kilobuck.
  • we should engage in more serious sensor design before Mariner, regardless of whether we take any action now.
    • as a starting point, the Lakeshore catalog and appendices (ref. product info page) and other resources should be absorbed, for considerations like thermal anchoring, lead length, benefits of 4-lead wiring, polyimide leads causing less conduction to the sensor than teflon, phosphor bronze having lower thermal conduction than copper, etc. Most of these topics are gathered from Appendix C.
Attachment 1: RTD_Resistance_vs_Temperature_for_PPG102A6.png
  2589   Wed Jun 16 17:17:12 2021 KojiUpdateGeneralI2 cell

I was searching an I2 (Iodine) cells back to the days of the laser gyro.

I found a likely box at a very tricky location. Took the photos and returned to this tricky place.

2021/Jul The box was moved to the OMC lab (KA)

Attachment 1: P_20210616_170104.jpeg
Attachment 2: P_20210616_170038.jpeg
Attachment 3: P_20210616_170021.jpeg
  2588   Fri Jun 11 16:48:31 2021 Aidan, StephenSummary2um PhotodiodesChamber is leaking

We hit another dead-end with leak hunting the IR labs dewer (we replaced screws and helicoil on the valve connection but there is still a big leak). We cleaned the flange and O-ring with isopropanal and replaced the threads with helicoil but still get the same sort of leak where we only hit 1E-2 Torr after 5 minutes of pumping and stablize around 1E-3.

After turning off the pumping station, the pressure rose quickly to 1Torr (in roughly 10 minutes or so).



[Stephen, Aidan, Wednesday 09 June]

Summary and Plan:

  • Poor sealing at the output valve (The Gap) needs to be resolved.
  • Planning to install #4-40 helicoils today (chamber will remain sealed, will need to remove output valve and cover output orifice, then transport the chamber to the WB EE shop for redrilling of holes.)
  • Meeting with Nina and Aidan this afternoon to iterate one more time.

Troubleshooting steps taken:

  1. Aidan took us through the full sequence of pump down and disassembly to bring me up to speed.
  2. We opened the lid and inspected the old o-ring.
    1. Signs of plastic deformation and of small flecks of particulate near sealing surface - good idea to change.
  3. We found new o-rings in a box from the Cryo lab, and one of these was swapped in after a good wipedown with IPA.
  4. Upon pumpdown, Aidan compared behavior and found no meaningful change to rate of pumpdown or stable pressure in e-4 torr range after 10+ minutes.
    1. By valving off [chamber + gauge] from pump line, it was clear that there was a leak in that volume, as within seconds the pressure rose from e-4 torr to e-2 torr, and stabilized at _(need to confirm - e0?)_ torr over ~10 minutes.
  5. Attempted to squirt IPA along o-ring seals, but there was not good access to the sealing surfaces, so this was a null test
  6. Looked closer at all of the chamber features, and noticed The Gap between chamber wall and chamber output valve, pictured in [Attachment 1]. Not good! But promising as a leak source.
    1. Three of the four screws were found to be loose due to apparent thread damage.
  7. IPA was squirted into The Gap at stable pressure of e-4 torr, but no change in pressure was noticed.
  8. Longer #4-40 screw reinstallation was attempted, and I could feel a small amount of pull at the very tip of the screw, but tightening the screws led to that small pull to fail as well - need to rework.
  9. The Clamp was installed and The Gap was closed [Attachment 2].
    1. When isolated from the vacuum pump, the chamber pressure progressed more slowly. Within seconds, we were at e-3 torr, and over 10 minutes the pressure stabled at e-2 torr, about 30x lower pressure per Aidan's records.


Attachment 1: IMG_3002.jpg
Attachment 2: IMG_3001.jpg
  2587   Fri Jun 11 16:00:31 2021 RadhikaDailyProgressCryo vacuum chamberLarge chamber heating

Summarizing heater trends from Wednesday, 6/9. Reiterating from post [2584], I ran the CTC-100 temp controller's auto tune routine to adjust PID coefficients for the heater. I then set the setpoint to 123K (operating temperature); the controller takes in the workpiece temperature as feedback. The heating data for this period is attached.

It took under 30 minutes for the temperature to rise from 72K (current cooling limit) to within a degree of the setpoint, 123K. The power delivered to the heater (bottom plot of first attachment) stayed below 25W, the limit we hardcoded. The workpiece temperature rises smoothly and plateaus around the setpoint, without significant overshooting. The controller holds the setpoint temperature pretty well thereafter. The second attachment is zoomed in on the workpiece temperature alone.

This run served as a test of the temperature controller's stability at the desired setpoint. Moving forward, we will continue to improve the cooling capacity of the chamber guided by our model. Once we optimize how cold we can get, we now know the heater can hold us at the desired temperature setpoint.

Attachment 1: heating_6-9.pdf
Attachment 2: heating_6-9_workpiece.pdf
  2586   Fri Jun 11 07:42:54 2021 StephenSummary2um PhotodiodesChamber is leaking

[Stephen, Aidan, Wednesday 09 June]

Summary and Plan:

  • Poor sealing at the output valve (The Gap) needs to be resolved.
  • Planning to install #4-40 helicoils today (chamber will remain sealed, will need to remove output valve and cover output orifice, then transport the chamber to the WB EE shop for redrilling of holes.)
  • Meeting with Nina and Aidan this afternoon to iterate one more time.

Troubleshooting steps taken:

  1. Aidan took us through the full sequence of pump down and disassembly to bring me up to speed.
  2. We opened the lid and inspected the old o-ring.
    1. Signs of plastic deformation and of small flecks of particulate near sealing surface - good idea to change.
  3. We found new o-rings in a box from the Cryo lab, and one of these was swapped in after a good wipedown with IPA.
  4. Upon pumpdown, Aidan compared behavior and found no meaningful change to rate of pumpdown or stable pressure in e-4 torr range after 10+ minutes.
    1. By valving off [chamber + gauge] from pump line, it was clear that there was a leak in that volume, as within seconds the pressure rose from e-4 torr to e-2 torr, and stabilized at _(need to confirm - e0?)_ torr over ~10 minutes.
  5. Attempted to squirt IPA along o-ring seals, but there was not good access to the sealing surfaces, so this was a null test
  6. Looked closer at all of the chamber features, and noticed The Gap between chamber wall and chamber output valve, pictured in [Attachment 1]. Not good! But promising as a leak source.
    1. Three of the four screws were found to be loose due to apparent thread damage.
  7. IPA was squirted into The Gap at stable pressure of e-4 torr, but no change in pressure was noticed.
  8. Longer #4-40 screw reinstallation was attempted, and I could feel a small amount of pull at the very tip of the screw, but tightening the screws led to that small pull to fail as well - need to rework.
  9. The Clamp was installed and The Gap was closed [Attachment 2].
    1. When isolated from the vacuum pump, the chamber pressure progressed more slowly. Within seconds, we were at e-3 torr, and over 10 minutes the pressure stabled at e-2 torr, about 30x lower pressure per Aidan's records.
Attachment 1: IMG_8865.JPG
Attachment 2: IMG_8881.JPG
  2585   Thu Jun 10 15:13:26 2021 RadhikaDailyProgressCryo vacuum chamberCooldown modeling vs. data

Code for simple heat transfer modeling can be found here: https://git.ligo.org/voyager/mariner40/-/blob/master/CryoEngineering/qil_simple_heat_transfer.ipynb

My original 1D cooldown script modeled conductive cooling along the copper braid as: Pcool = k * A/L * (T - T_set); where Pcool is the cooling power, k is the thermal conductivity of copper, A is the cross-sectional area of the braid, and L is the length of the braid. I changed the code to instead use the tabulated power vs. temperature points for the CH-104 coldhead, taken from Paco's script qil_heat_estimate.ipynb. The first figure compares the interpolated curves from the tabulated values (at 50Hz and 60Hz operation), to the original conductive transfer model (50K setpoint). The original conductive power-temp relationship is linear, which overestimates the cooling power at high temperatures. Switching to the tabulated points results in more realistic model. Moving forward, I intend to use the 50Hz interpolated curve. 

The script considers radiative heating to the coldplate from the the chamber bottom (rough aluminum) and the outer shield (coated in aluminum foil). It assumes over a long period of time that the inner shield and coldplate temperatures are equal.

The second figure shows the results of this model alongside the actual coolddown data extracted from the CTC-100. It is clear that the model is not accounting for additional radiative heat sources that would explain the slower cooldown and higher final temperature. Adding in model complexity is my current focus. 


Attachment 1: coldhead_capacity.pdf
Attachment 2: cooldown_model_vs_data.pdf
  2584   Wed Jun 9 13:35:07 2021 RadhikaDailyProgressCryo vacuum chamberHeater actuation

Today I attempted to auto tune PID coefficients for the heater, so that we can reach and maintain a setpoint of 123K with appropriate ramp-up. The workpiece was around 72K originally. For auto tuning, I set the max power for the heater to 25W. I adjusted the lag time to 30s, and changed the setpoint to 72K so that the tuning response measured how stable the system is when being perturbed from our set point. The auto tune process ran without an error; however, by default the tuning mode switched to "step" tuning, and I am not sure why this occurred. The tuning took roughly 5 minutes to complete. The final message is attached; the adjusted parameters were:

Gain: 2.703

Lag: 125.8 s

Time constant: 271.9 s

I was expecting the adjusted output to be new PID coefficients, but I noticed that the PID coefficients did seem to change after this process. To begin warmup to 123K, I changed the setpoint to 123K and let the heater do its thing (the feedback temperature is set to that of the workpiece). The heater power stabilized to around 17W, and the temperature of the workpiece reached within a degree K of the setpoint within 30 minutes. I am letting the temperature hold at 123K overnight and plan to return tomorrow morning to check on it and extract the heating data for plotting. 

Attachment 1: IMG_1722.jpeg
  2583   Tue Jun 8 12:58:40 2021 StephenDailyProgressCryo vacuum chamberMaximum power for temperature actuation

Quick log establishing the maximum power for our thermal actuation:

Heater: HSA25100RJ from TE, unknown sourcing. Acetone wiping cleaned off p/n and markings from body, should engrave at next opportunity, but [Attachment 1] from many months ago shows the p/n. Note that this is not the current mounting configuration - [Attachment 2] is more similar to current mounting. Anyway, according to the datasheet (now added to the QIL wiki at Documentation > Manuals) this heater is rated for 25W and has a resistance of 100Ω.

Leads: unknown, and not super important unless we had tiny hair conductor - I am not in lab presently, but it appears from our connector (Lesker FTACIR19AC) that we must have 20-24 AWG,

Carrying the 7, the current through the Heater will be 0.25 A at max actuation, and the 20-24 AWG insulated copper leads will have plenty of ampacity for this load (plus, they are cooled, so normal current capacity considerations fly out the window a bit).

Conclusion: 25W actuation will be the limit that we will apply in the CTC100 temperature actuation routine.

From first trials yesterday, the response at ~20W (at starting temperatures around 80K) appears to be on the order of 1 degree per minute, which should be just fine for actuating to maintain a +/-  1 degree constant setpoint with static thermal loads. More to follow on trials implementing temperature control.

Note that the QIL Wiki points to the DCC (which contains a budget that was helpful resources to trace these purchases), the datasheets and other documentation, and also points to the QIL Cryo Vacuum Chamber photo album, which hosts the images below.

Attachment 1: IMG_8495.JPG
Attachment 2: IMG_8745.JPG
  2582   Fri Jun 4 16:15:00 2021 RadhikaDailyProgressCryo vacuum chamberCooldown from 6/2-6/4

I extracted cooldown data from the CTC100 USB around 1pm today (~50 hours of cooldown). I've attached a log plot below. The heater RTD seemed to be behaving weirdly at the beginning, but soon stabilized and cooled as expected.

I estimated the time constant for the workpiece: 50 hr/ (300K - 90K) = 0.24 hr/K =  ~860 s/K



Attachment 1: cooldown_6-2_6-4.png
  2581   Wed Jun 2 20:06:07 2021 ranaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

I'll be curious to see the results of Radhika's thermal model - I am suspicious of this thermal strap contact to the base plate. It would be good if we could instead make a copper mating plate:

  1. groovy milled out loop holder for the strap to go into such that the contact surface is larger
  2. flat base to get smushed into the colde plate
  3. thru holes so that it can be directly screwed down with some washers
  4. screwed down with a torque wrench so that we know that its really tight (i.e. "I think its pretty tight" doesn't really work when we want a good thermal contact)
  5. someone with a MechE degree calculates the proper torque such that we plasticly deform the copper and get it to sit more flush to the cold plate
  6. a cap for the copper plate that really smooshes the copper strap into the bottom part of the mating plate.
  2580   Mon May 31 11:38:13 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

[Stephen, Radhika]

After Thursday's work, we resumed on Friday and lifted up the coldplate to access the collar and baseplate. Stephen added metallized mylar wraps to the peek cylindrical spacers [pic 9, 10]. He took out the cylindrical collar and baseplate (previously there to minimize contact between the colplate and chamber bottom) and replaced them with a foil collar [pic 13, note this is before pushing down the foil over the PEEK spacers]. We re-inserted the coldplate.

In order to wrap/insulate the copper braid, we inserted 2 sheets of metallized mylar into the vacuum tube to surround the braid [pic 11, 14]. We cut the mylar wrapping so that it did not short to the outer shield. We also confirmed that the copper braid is not shorting to the inner shield hole (there was clearance for the shield to be lifted before hitting the braid). The mylar extends to the coldhead, which we then wrapped with aluminum foil, making sure not to short to the walls of the tube [pic 12]. 

We switched the foil wrapping from the inside to the outside of the outer shield, so that any radiative transfer from the inner shield to the outer shield would be absorbed as much as possible (not reflected back). We placed both shields back and bolted down the copper braid loop and workpiece. We re-attached the RTDs, then placed in both shield lids (without bolting them down) and stopped for the day.

Today I checked on the chamber setup and closed up. I started the vacuum pump and it made abnormally loud noises, indicating something was off. The percentage sign continued to flash, indicating that the pump was not reaching 80% speed. I performed 2-3 power cycles, which did not solve the issue. We will pick up to debug the issue early this week.

[update 02 June 2021]

Pumpdown and cooldown were successfully started this morning - Radhika retightened the green leak valve and pumps started just fine. We will check trends on Friday and likely will allow cooldown to proceed over the weekend.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album.

Attachment 1: IMG_1681.jpeg
Attachment 2: IMG_1684.jpeg
Attachment 3: IMG_1686.jpeg
Attachment 4: IMG_1687.jpeg
Attachment 5: IMG_1688.jpeg
Attachment 6: IMG_1690.jpeg
Attachment 7: IMG_1693.jpeg
Attachment 8: IMG_1695.jpeg
Attachment 9: IMG_8796.JPG
Attachment 10: IMG_8805.JPG
Attachment 11: IMG_8811.JPG
Attachment 12: IMG_8815.JPG
Attachment 13: IMG_8803.JPG
Attachment 14: IMG_8810.JPG
  2579   Thu May 27 16:48:35 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling --> started warmup

This morning I vented and opened up the chamber to add aluminum foil and other insulation. Stephen's order of mylar sheets, peet sheets, and G10 rings came in today and I picked it up from Downs.

⁃ unscrewed the bottom conflat of the tee and added foil where I could reach (junction of cold head and copper braid) [pic1]

- cut narrow rings of 0.01'' thickness peet sheet and inserted into the main hole of the inner/outer shields, to prevent shorting to the copper braid [pic 6, 2, 3]

- added foil to the inside of the outershield and outside of the inner shield, poking holes for all viewports [pics 4,5]

⁃ unclamped the copper braid from the coldplate and the RTDs from all locations (shields, heater, workpiece mount) to prepare for lining baseplate with foil [pic 7]

Tomorrow we plan to wrap the shaft of the copper braid with metallized mylar and an additional layer of foil. I left both shields outside of the chamber so that tomorrow we are ready to remove the cold plate and add foil below. 


  • Cooldown complete as of 4:41, with part 2 of data posted to CSVlogs Box folder with file name "cooldown part 2 20210525.CSV" for Radhika's analysis.
  • Resumed USB logging and started Warmup. Simply turned off cryocooler using compressor panel's green power switch, then flipped circuit breaker lever to off position.
    • Expected to be ready for opening, with warmer-than-freezing temperatures throughout, on Thursday morning.
    • Noticed a "burp" of the gauge pressure readings up to e-2 torr range immediately upon turning off cryocooler, lasting about 1 minute before gauge readings were back down to e-6 range. Interesting, and something to keep an eye on.
  • Linked Active Ion Gauge (5e‐2 to 5e‐10 torr per AIG manual) to Active Pirani Gauge (atm to 7e‐3 torr per APG manual) so that the Edwards TIC controller will enable the high-vacuum AIG during pump down and disabled during venting. Prior, the AIG had to be manually enabled at suitable operating pressures. Followed instructions at Section 4.11 of the TIC controller manual. Recall that all manuals are hosted on the QIL Cryo Vacuum Chamber wiki.

Radhika, feel welcome to post full cooling data to this entry, or to your original - up to you!


Attachment 1: IMG_1678.jpeg
Attachment 2: Photo_on_5-27-21_at_15.53.jpg
Attachment 3: Photo_on_5-27-21_at_16.01.jpg
Attachment 4: Photo_on_5-27-21_at_16.22.jpg
Attachment 5: Photo_on_5-27-21_at_16.35.jpg
Attachment 6: Photo_on_5-27-21_at_16.38.jpg
Attachment 7: Photo_on_5-27-21_at_16.37.jpg
  2578   Wed May 26 18:38:25 2021 AidanSummary2um PhotodiodesChamber is leaking

I tried Krytox around the O-ring and also tightening the screws around the valve. The leaking persists at roughly the same rate.


I tried pumping down the JPL PD chamber to test the new PD at cryo temperatures. Unfortunately, the chamber can;t get past about 6E-3 Torr with the pump on. As soon as I turned off the pump the pressure rose to around 2 Torr over 20 minutes or so.

I extricated the chamber from the pedestals, flipped it and removed the bottom plate. I cleaned the O-ring with isopropanol and wiped down the mating surface on the chamber (also with iso). I replaced the plate and tightened the screws. Then I returned the chamber to the table and reconnected it to the vacuum system. I tried pumping down once again but I saw pretty much exactly the same situation as before (pressure bottoming out around 6E-3 Torr and then rising quickly again when the pump was turned off).

I guess it's possible that the O-ring is damaged - although I couldn't see anything obivous. We didn't mess around with the viewport (when we replaced the diode a few weeks ago) so I'm hoping there is no issue there.


  2577   Tue May 25 16:38:17 2021 StephenDailyProgressCryo vacuum chamberWeekend cryo chamber cooling --> started warmup
  • Cooldown complete as of 4:41, with part 2 of data posted to CSVlogs Box folder with file name "cooldown part 2 20210525.CSV" for Radhika's analysis.
  • Resumed USB logging and started Warmup. Simply turned off cryocooler using compressor panel's green power switch, then flipped circuit breaker lever to off position.
    • Expected to be ready for opening, with warmer-than-freezing temperatures throughout, on Thursday morning.
    • Noticed a "burp" of the gauge pressure readings up to e-2 torr range immediately upon turning off cryocooler, lasting about 1 minute before gauge readings were back down to e-6 range. Interesting, and something to keep an eye on.
  • Linked Active Ion Gauge (5e‐2 to 5e‐10 torr per AIG manual) to Active Pirani Gauge (atm to 7e‐3 torr per APG manual) so that the Edwards TIC controller will enable the high-vacuum AIG during pump down and disabled during venting. Prior, the AIG had to be manually enabled at suitable operating pressures. Followed instructions at Section 4.11 of the TIC controller manual. Recall that all manuals are hosted on the QIL Cryo Vacuum Chamber wiki.

Radhika, feel welcome to post full cooling data to this entry, or to your original - up to you!

Attachment 1: cooling_full_20210525.png
ELOG V3.1.3-